The scale of two similar rectangles is 3:4. The area of the smaller rectangle is 207 square centimeters. What is the area of the larger rectangle?
Answer by edjones(8007)

Lw=A
Lw=207
Let L=w
w^2=207
w=sqrt(207)
=sqrt(9*23)
=3sqrt(23)
So, for the larger w=4sqrt(23)
w^2=(4sqrt(23))^2
=16*23
=368 cm^2 Area of the larger rectangle.
.
Ed